### v2

# The modelling of the action potentials in myelinated nerve fibres

Abstract
The classical Hodgkin-Huxley equations describe the propagation of an axon potential (AP) in unmyelinated axons. In many cases the axons have a myelin sheath and the experimental studies have then revealed significant changes in the velocity of APs. In this paper, a theoretical model is proposed describing the AP propagation in myelinated axons. As far as the velocity of an AP is affected, the basis of the model is taken after Lieberstein, who included the possible effect of inductance that might influence velocity, into the governing equation. The proposed model includes the structural properties of the myelin sheath: the -ratio (the ratio of the length of the myelin sheath and the Ranvier node) and g-ratio (the ratio of the inner-to-outer diameter of a myelinated axon) through parameter . It is demonstrated that the Lieberstein model can describe all the essential effects characteristic to the formation and propagation of an AP in an unmyelinated axon. Then a phenomenological model (a wave-type equation) for a myelinated axon is described including the influence of the structural properties of the myelin sheath and the radius of an axon. The numerical simulation using the physical variables demonstrates the changes in the velocity of an AP as well as the changes in its profile. These results match well the known effects from experimental studies.
keywords:
action potential, nerve fibre, myelinated axon, velocity, mathematical modelling1 Introduction
The celebrated Hodgkin-Huxley (HH) equations can describe the action potential (AP) in an unmyelinated axon taking into account sodium and potassium ion currents [1]. The strength of the HH equations is in the detailed description of the physics of ion currents but the cable equation describing the propagation of an AP is simplified by neglecting the inductance. There are several ways to improve this classical model towards a better description of the structural properties of axons and taking into account the accompanying effects. Many studies are devoted to the description of mechanical and/or thermal effects accompanying the propagation of an AP [2, 3, 4, 5] resulting in the formation of an ensemble of waves. These theoretical models are based on experimental results [6, 7, 8, 9, 10, 11]. Another important avenue of studies is related to modelling the behaviour of an AP in myelinated axons. The starting point for the modelling of an unmyelinated is a cylindrical tube embedded into the extracellular fluid. The barrier between extra- and intracellular fluids is a lipid bilayer with proteins. Such a biomembrane is sometimes modelled like a simple lipid bi-layer only [2]. In the case of the myelinated axon, this barrier has a myelin sheath which consists of multiple layers of a glial membrane composed of lipids and proteins and serves as an insulator [12, 13] or in some scenarios even as signal modulating element [14]. It means that under the myelin sheath the ion currents through the basic biomembrane (barrier) proposed by Hodgkin and Huxley [1], do not exist. However, the myelin sheath is interrupted by Ranvier nodes where the usual HH equations work. It is proposed that under the myelin sheath, the passive cable equation (the diffusion-type equation) describes the process and in Ranvier nodes, the usual HH equations work [15, 16]. However, it is demonstrated experimentally that neurons which have myelin sheathing often show higher AP velocities compared to similar unmyelinated neurons. This effect was already reported by Lillie in 1924 [17] and is nowadays referred to as the Lillie transition [18] or saltatory conduction (see [19]).
In this paper, a phenomenological model is proposed for describing the propagation of an AP in a myelinated axon. The most important question is how to model the possible changes in the propagation velocity of an AP due to the myelin sheath. The process is nonlinear and we follow a recommendation of Whitham [20]: “… one should not always turn too quickly to a search for the .” In the context of our equation, it means that we should keep the neglected inductance as small as it is (which we will see in Section 2.3 is about 40 [mH cm]). This idea is supported by Wang et al [21] who have argued that inductance is “a missing piece of neuroscience”. They state that the main source of the inductance is the myelin sheath. In general terms, however, inductance means that “there is a kind of biological structure that can store energy in a non-electrical form" [21]. If we keep the inductance in the cable equation then under the myelin sheath the propagation is described not by a diffusion (see [16]) equation but by a hyperbolic equation. This hypothesis may better explain the changes in the velocity of an AP. Given reports [22] noting the dependence of velocity on the diameter of a fibre, this relationship should be taken into account.
In what follows, the model of Lieberstein [23] is a basic one to describe the AP propagation in axons. This model includes the HH ion currents but in addition, includes the inductance as it follows from the Maxwell equations. For proper modelling of the processes in myelinated axons, the influence of the myelin sheath must certainly be taken into account. In principle, the length of a myelinated section and its thickness are the leading structural factors in the process. Consequently, one important parameter is the ratio of lengths of a node of Ranvier and a myelinated section in the form of the myelin length ratio (or the -ratio for short) which is proportional to . Another important parameter is the g-ratio (see [22]) which is the ratio of the inner-to-outer diameter of a myelinated axon. These parameters take into account the structure of a myelin sheath (c.f. analysis by Basser [24]).
The main idea of this study is based on the following considerations. First, when dealing with a nonlinear system, we follow Whitham’s idea [20] that all possible small terms (influences) should be taken into account. Second, in constructing a mathematical model for the propagation of an AP in myelinated axons, we start from Maxwell equations [25]. Third, for the description of structured properties of the axon, we use the phenomenological approach following the ideas of Hodgkin and Huxley [1] for describing the ion currents and Bressloff [26] for describing the saltatory conduction by a coupling coefficient which characterises the discrete diffusion. The general principles of phenomenological modelling are described by Engelbrecht et al [27].
Section 2 is devoted to the analysis of the Lieberstein [23] model for the unmyelinated axon. The results confirm the earlier analysis by Kaplan and Trujillo [28] but more details are given. As novel results, the profiles of ion currents and changes of the phenomenological variables , , and [1] during the propagation of the AP are also presented. It is shown that the head-on-collision of two APs leads to annihilation. The influence of the refractory period for the formation of consecutive APs is also analysed as well as the dependence of the velocity on the diameter of the axon. These well-known effects demonstrate the validity of the Lieberstein equation and encourage us to use this equation for the modelling of the APs in myelinated axons. Section 3 describes experimental results on myelinated axons focused on velocity changes [29, 30, 31], This analysis permits us to propose in Section 4 a phenomenological model describing the propagation of an AP in a myelinated axon influenced strongly by the structure of a myelin sheath (-ratio). The structure of the governing equations follows the Lieberstein model (see Section 2) where the inductance leads to a wave-like behaviour but the final velocity of an AP depends on ion currents. The numerical simulations demonstrate clearly the changes in the velocity under the myelin sheath of the propagating AP. Based on these results, a short discussion of forms of energy in an AP is presented in Section 4.2. Finally, in Section 5 the discussion is presented with conclusions. As far as the numerical simulations are carried on with physical variables (i.e., in physiologically viable range), the results could be better checked in experimental studies.
2 Lieberstein model for AP propagation on unmyelinated axon
Experimental studies of processes in the myelinated axons have demonstrated that the signal propagates faster than in unmyelinated axons [32, 33, 34]. Note that Cohen et al. [35] argue about elusiveness of the role of submyelin conduction. Although there are several explanations of this phenomenon [36, 26, 15] there is a need for constructing a mathematical model which allows to calculate of an AP profile in a myelinated axon with the velocities depending on structural properties of the myelin sheath. Consequently, in deriving a mathematical model, all possible phenomena that could influence the propagation velocity should be taken into account, even if some of these could be considered much smaller than others at first glance (see, for example, Whitham [20]).
2.1 Basic model
The cable equation used by Hodgkin and Huxley [1] for deriving the equations for the electrical signal propagation is parabolic, i.e., inductance is neglected. Indeed, its influence is small but following Whitham [20] we use here the full model of Lieberstein [23] derived directly from Maxwell equations [25]. This model is hyperbolic but the final velocity of the signal is influenced by ion currents like in the HH equations [1].
Here we go briefly over the key points from [23] demonstrate the importance of a threshold and annihilation and present the profiles of ion currents and phenomenological variables , , (not presented in [23]). We also add some comments to make it easier to follow later in the paper what is changed and why is changed when including the effects of myelination in the next Section.
Lieberstein [23] starts with an elementary form of Maxwell equations for current and voltage on a long line
where
is space (length) and is time;
is the action potential, is the line axon current (along the axon) and is the membrane current per unit length (taken the same as HH current across the membrane later in the paper);
is the radius of the axon, is the axon resistance per unit length, is the axon specific self-inductance and is the axon self capacitance per unit area per unit length.
It is noted that the membrane current density (which is what is used in the HH equations [1]) is
and the specific resistance of the axon is
The structure and the physical interpretation of the governing equations (1) and (2) is easier to see if writing it as a single second-order PDE (see eq. (3) in [23]):
where on the left-hand side (LHS) there is a classical wave equation type operator while on the right-hand side (RHS) there are dissipation-type operators (first order partial time derivatives) which either remove energy from the system or add it to the system depending on the sign of the operator. In the context of signal propagation in nerve fibres, the ionic currents across the membrane change the membrane potential in a given location while some of that energy is lost during the diffusion-type propagation along the axon to the axon resistance. At the same time, the LHS operator ensures that part of the signal is propagating like a classical wave.
Returning to the system as two coupled eqs. (1) and (2) the membrane current is taken (this is what is different from regular telegraph equation) as it was proposed in the HH equations [1] as
where is the total membrane current density (positive direction is into the axon), is the membrane potential relative to the resting potential (depolarisation is negative), is the membrane capacity per unit area while and are currents which are defined through “equilibrium potentials" for Na and K ions while is “leakage current" representing all other ionic currents (Hodgkin and Huxley point mostly at chlorine Cl, however, Ca is also potentially relevant which some authors claim to be about as important as Na and K ions) and is chosen so that at resting potential leakage current across the membrane would be zero.
In eq. (1) we have membrane current per unit length , which is related to the membrane current density as and if one takes as expressed in eq. (6) inserting it into eq. (1), we can get expression
Further, inserting [1] and collecting both time derivatives in eq. (7), we get:
and from eq. (2):
The dimensions of coefficients in eqs. (8) and (9) are: , , , are dimensionless, ( denotes SI-unit: [Siemens]) while the membrane potential and the line axon current . As usual, the space dimension is , time dimension is , while in addition dimension is and dimension is .
The difference in keeping the inductance contrary to the celebrated classical HH equations [1] is the following. One possible way of making sense of the differences is by looking at the equations. In the case of the system of PDEs (eqs. (1) and (2) in [23]), it can be seen that there is a term in eq. (2). This can be interpreted as the rate of change of current along an axon. In the HH equations, the axon current is “hidden” in term . Another possibility is to look at the Lieberstein model in the form of one PDE (eqs. (3) and (6) in [23]). It can be seen in eq. (3) that two additional terms appear in the Lieberstein model compared to the classical HH equations – and . The second partial derivative is what makes the model hyperbolic and as pointed out by Kaplan and Trujillo [28], it influences the maximum velocity of the AP. However, the term is an even more notable difference as the time derivative of the ion current is not present in the classical HH equations. Observing eq. 6 in [28] it can be seen that inductance affects the ion currents and also time derivatives of gating variables arise. One should note that Kaplan and Trujillo [28] have calculated axon inductance from the movement of ions as if AP velocity is at µ and Cole and Baker have also investigated question of inductance in this context [37, 38]. It can be noted that similar models involving inductivity are in use in cardiophysiology, see, for example, [39].
2.2 Moving frame of reference and inductivity
Lieberstein [23] assumes that accounting for the inductivity (unlike in the HH equations where it was neglected) he can go into a moving frame of reference in a standard way introducing However, we should note that it could be, actually , where is velocity, as taking only a single direction means discarding the wave propagating in the opposite direction. It should be noted that using a moving frame of reference allows one to also derive an evolution equation (a single wave) for the nerve pulse [40]. Lieberstein [23] defined the velocity of the AP in an unmyelinated axon through axon radius , inductance and capacity as
proceeding after that to writing up eqs (1) and (2) in the moving frame of reference. For the sake of completeness, it should be noted that in (10) Lieberstein [23] argued that for capacitance the axon self-capacitance contribution can be neglected because and axon radius is small so the contribution of is small compared to membrane capacity per unit area .
More recently, Fukasawa and Takizawa [41] have also investigated the question of electrical signal propagation velocity in an axon if inductivity is taken into account including the question of how to estimate the inductance parameter value.
In the present paper, we have opted to avoid going into a moving frame of reference preferring to solve the model as a coupled pair of PDEs and to keep the influence of for the sake of completeness, even if its influence on the behaviour of the solutions is arguably small.
2.3 Solutions of Lieberstein model
A numerical example (see Fig. 1) using pseudospectral method (PSM) (see Appendix A in [2]) and periodic boundary conditions demonstrating the evolution of solutions for model equations (8) and (9) is now presented. Briefly, the main point of the pseudospectral method is that the discrete Fourier transform (DFT) based (PSM) (see also [42]) can be used to represent variable in the Fourier space as
where is the number of space-grid points, is the space step, ; is the imaginary unit, denotes the DFT and denotes the inverse DFT. The idea of the PSM is to approximate space derivatives by making use of the DFT
reducing, therefore, the partial differential equation (PDE) to an ordinary differential equation (ODE) and then to use standard ODE solvers for integration in time. For integration in time, the model equations are rewritten as a system of first-order ODE’s and a standard numerical integrator is applied. In the numerical examples given in the present paper the ODEPACK FORTRAN code (see [43]) ODE solver is used through its NumPy implementation. Handling of the data and initialization of the variables is done in Python by making use of the package SciPy (see [44]) and the numerical results are analysed and visualised in the Matlab environment.
The following parameters are used: (number of spatial nodes), (time in [ms]), µ (membrane capacitance), µ (axon radius), (axoplasm resistance) while the HH equations parameters are the same as taken in [1]: and while (initial values for parameters in HH equations at ). We take in the following example the axoplasm capacitance as µ and (resistance of an axon per unit length). We take (inductivity) for the numerical example (resulting in 1 propagation velocity for the AP signal). We remark that, as noted earlier by Kaplan and Trujillo [28], have calculated axon inductance from the movement of ions as if AP velocity is at µ. However, as in our case the reference axon radius is taken smaller it makes sense to reduce the value. Initially system is taken at rest and at we generate a narrow bell-shaped pulse (“spark") for the in the middle of the space domain which generates the propagating AP. In spatial units, the length of the computation node () is 61 while the width of the computational domain in the space (from to ) is 4. For the sake of readability, the additional equations and values of physical quantities used in numerical simulation
are collected in Table 1 following [1] and [23]:
| Table 1: Parameters for finding example solutions for (8) and (9). | ||
It should be stressed, that the chosen parameters do not represent any particular model nerve as these have been collected from studies describing different experimental setups or even taken as a rough estimate (like specific inductivity L). Most of the parameters are taken from [1] corresponding to giant squid axon at about while inductivity is chosen so that the AP signal would have propagation velocity of around 1 if the axon radius is 1 [µ. It can be noted that the used value of inductance is large (40 [mHcm]) compared to, for example, a solid copper wire which would have inductance of roughly 18 if the radius of the wire is 1 µ and the length is 1 . Although, clearly, the structure of the axon is very different from a solid metal fibre, the difference of roughly 6 orders of magnitude is significant. What inductivity is, in principle, is a process that resists the change of magnetic field and considering various geometrical structures (like cytoskeleton) and different physical processes interacting with each other the chosen value does not appear to be outright implausible. One can note, for example, the significant mass difference between electrons and considered ions. However, it is clear that the nature of inductivity in the context of nerve fibres needs further clarification and study. We have a process in the proposed model that behaves like inductivity, and has a dimension related to inductivity so we have to make an assumption that it is inductivity to avoid violating Occam’s razor principle – although the relatively large value of the parameter needed to explain the experimental observations hints that there might be something more going on (see also Wang et al. [21]).
In Fig. 1 it is demonstrated that the Lieberstein model has solutions which are characteristic of an AP in earlier studies (for example, [1, 23]), however, it is known that in physiology, an AP has a few more properties that must be fulfilled. These are: (1) electrical nerve signal must annihilate upon head-on collision, (2) a disturbance or signal generated during a short time after the signal has passed a location (so-called refractory period) must dissipate rapidly without generating a new AP (a minimum separation time before another nerve pulse can propagate). In Figs. 2 and 3 it is demonstrated that the Lieberstein model satisfies these conditions. As we are using periodic boundary conditions (nodes and are connected so a wave hitting the boundary enters the grid again from , as an example). Figure 2 demonstrates that during a head-on collision of propagating AP signals annihilate according to the Lieberstein model fulfilling the first condition for a realistic AP model. The left panel of Fig. 2 depicts the AP amplitude in time at boundary node and the right panel depicts the spatial distribution of the signal when two spatial periods are plotted side by side (from to ) making it clearer to see a head-on collision of propagating AP signals at the right boundary. The left panel of Fig. 3 demonstrates the emergence of the AP from the initial disturbance (centred at ) above the threshold value given at and then later another voltage spike is given at the same location at which dissipates rapidly. The middle panel of Fig. 3 demonstrates that if a sufficient temporal separation between the pulses is given a repeating signal can be generated and the right panel of Fig. 3 shows a niche scenario where giving a small signal which is initially below the threshold value sufficiently frequently it is still possible to generate propagating AP if the combined dynamics succeeds in crossing the threshold value for the AP generation.
2.4 Influence of the axon radius on the propagation velocity of AP.
From earlier studies (see, for example, [45]) it is known that the propagation velocity of the AP is influenced by three factors – (i) space coefficient , (ii) time coefficient , and (iii) the time measure characterising the time it takes to generate the AP on any point along the axon. The membrane resistance scales inverse-proportionally with axon diameter while membrane capacitance increases proportionally with axon diameter while axial resistance scales as meaning that the propagation velocity of AP is .
For a start, a so-called “sanity test" was carried out using the data of the HH celebrated experiment [1] In this case the temperature was 6.3 [Co] and we chose [mHcm]. Then the simulation yields an AP propagation velocity of 19.0 [m/s] compared to 18.8 [m/s] in the HH paper [1] against the experimentally measured AP velocity of 21.2 [m/s].
We note that in the model parameters, the axon resistance is changed when the axon radius is changing so , where is constant [1]. The propagation velocity of AP as a function of axon radius is depicted in Fig. 4 with the parameters used for a numerical example. It should be noted that the electrical circuit hypothesis for the AP propagation velocity (like for the HH equations [1] or for the Lieberstein model [23]) which is based on similarity with the chosen electrical circuit, while certainly useful, can overlook some potentially important mechanisms (like, for example, the influence of cytoskeleton) that are left aside during the many simplifications needed to get a simple electrical scheme. This means that in practical applications one has to consider carefully any conflicts between experimental observations and modelling results to determine if the model is sufficient for adequately describing the observed process. For example, here we have chosen so that at [µ] the AP velocity is as a demonstration of the model behaviour, for practical applications that parameter would need to be determined from the electrophysiology of the problem under investigation as in reality, this is not a free parameter.
| Table 2: AP velocity as a function of axon radius . | ||||||||
| a µ | 0.25 | 0.50 | 1 | 2 | 4 | 8 | 16 | 25 |
| 0.5 | 0.75 | 1.0 | 1.4 | 1.95 | 2.55 | 3.2 | 3.55 | |
| a µ | 40 | 50 | 60 | 80 | 100 | 125 | 150 | 175 |
| 3.9 | 4.0 | 4.1 | 4.25 | 4.3 | 4.35 | 4.4 | 4.4 | |
3 Description of a myelinated axon
The axon can be described as a tube filled with axoplasm and cytoskeleton [13]. The barrier between extra- and intracellular fluid has a lipid-bilayer base but also has membrane proteins [46]. While some axons are unmyelinated, many of them have a myelin sheath (additional structure surrounding the axon formed by glial cells (oligodendrocytes (in the central nervous system) and Schwann cells (in the peripheral nervous system)) that is in simplified terms composed of layers of biomembrane glued together with some proteins.
3.1 The structure of myelinated axon
The myelin sheath is interrupted by the nodes of Ranvier which play an important role in the nerve pulse propagation. A node of Ranvier is typically around 1 µm in length and has a high density of ion channels [29] while myelin sheath segments are typically from roughly 50 µm to 300 µm in length. The distribution of the myelinated parts was once believed to be uniform, but now it is understood that the length of the myelinated parts and the nodes of Ranvier vary and unusually long nodes of Ranvier (50 µm) have been reported [12]. These long segments could play an important role in the synchronisation of nerve pulses by delaying an AP [30].
The myelin sheath is a stack of specialised plasma membrane sheets produced by glial cells that wrap helically around the axon [29]. A lipid bilayer (biomembrane) is typically from 3 to 4 nm in thickness and contains various proteins in and between the layers [47] that all together form a myelin sheath composed of many such layers wrapped helically around the axon. The myelin sheath can be up to 2.5 µm in thickness [48, 49, 50] and on the lowest theoretical limit, an extra layer of biomembrane surrounding the axon might be argued to be a myelin sheath. Typical axon diameter varies from about 0.5 µm to 20 µm [48] in mammals but can reach as high as about 1 mm in squid giant axon [51, 10]. For example, in peripheral nerve fibres, the myelin sheath thickness starts from about 0.5 µ for the smaller diameter axons and the upper value of 2.5 µ is more characteristic to the axons with large diameter [48]. In peripheral nerves, myelin sheath thickness increases sharply, at first, when the axon radius increases from the smallest physiologically viable diameters and then the thickness increase gradually slows down as the axon diameter approaches the largest physiologically viable values [48] in these nerves.
Nodes of Ranvier contain much higher densities of various ion channels than elsewhere on axons. K and Ca channels are about 10 nm of the radial length across the membrane and roughly 4 nm of the longitudinal diameter in the plane of the membrane [52]. Na channel is about 12 nm of the radial length across the membrane and about 10 nm of the longitudinal diameter in the plane of the membrane [53]. Structurally the myelinated part of the axon is divided into the following regions: next to the node of Ranvier is a region called the paranode. This is the area where the myelin attaches to the axon. The juxtaparanode is located next to the paranode and it is the area where most voltage-gated K+ ion channels are located. The Na+ channels are concentrated in the nodes of Ranvier. The geometry of a myelinated axon is sketched later in Fig. 5. A simplified model of the sheath can be presented as a stack of lipid bilayers. It should be noted that in the proposed model we do not consider different ion channel distributions between juxtaparanode and paranode and take uniform ion channel distribution within the node of Ranvier. The AP propagates down an axon without diminution at velocities up to 1 m/s (without myelin sheath) [29]. In non-myelinated neurons, the conduction velocity of an action potential is roughly proportional to the diameter of the axon [29]. The presence of a myelin sheath around an axon increases the velocity of impulse conduction to 10-100 m/s [29, 30]. This is a brief overview of the structure of the myelinated axons needed for further modelling. The detailed overview of the properties of myelinated nerve fibres is presented in [34, 54].
3.2 Saltatory conduction mechanism hypothesis for myelinated axon
It is known that the myelination of the axon increases the propagation velocity of the AP and the prevalent explanation in earlier studies of this phenomenon is the saltatory conduction mechanism [26, 15, 55, 19, 32]. As the capacity of the axon changes significantly between the myelinated and unmyelinated sections, this causes, in a nutshell, the electrical signal to “jump" between the nodes of Ranvier propagating faster than it would in the case of an unmyelinated axon. The saltatory conduction hypothesis is briefly summarised as follows.
In earlier studies, it is demonstrated that myelination increases effective membrane resistance (reduces the permeability of ions) and decreases the capacitance of the membrane by several orders of magnitude. As noted, the propagation of AP along myelinated axons is considerably faster than in unmyelinated axons. In the context of the cable equation (which is a starting point for both HH and Lieberstein models) the usual explanation in literature is that transmembrane currents in the myelinated sections can be neglected meaning that the myelin sheath section can be taken in that case as a simple resistor (i.e, the sheath acts as an insulator). This phenomenon is called a saltatory (leaping) wave in earlier studies where AP is not propagating continuously along the axon but rather jumps from one node of Ranvier to another node of Ranvier [26].
In practice, the saltatory conduction hypothesis, although difficult to verify experimentally as measuring sharply localised AP (nodes of Raniver are typically around 1 µ long) is measured [19]. What is measured in classical experiments (for example, [6, 56, 51]), is AP over some section of the axon (i.e, including the signal from both myelinated and non-myelinated sections, moreover most classical experiments are done on non-myelinated axons). However, the proposed hypothesis appears to be a plausible (and widely used in earlier studies) starting point to explain why the AP is faster in myelinated axons and we opt to use it as well. It should be noted that in effect we use interpretation which is partly inspired from the continuum mechanics [57] where we interpret the signal coming from the model not as the evolution of AP in time at a fixed point on axon as is typical but rather a combined signal one might get from a “unit cell" which contains a node of Ranvier and a myelinated section next to the node, based on assumption that characteristic wavelengths of the signal are much larger than the underlying scale of the micro-structure (nodes of Ranvier in µ range and myelinated sections typically in hundreds of µ in length [29, 13]). In comparison, the first harmonics (when looking at the Fourier spectrum of the signal) of the AP in space are from to tens of (depending on the duration and velocity of the signal) [58, 59].
Following [26], it is assumed that membrane potential is uniform within a given node of Ranvier (i.e, a node is an isopotential) and denoting the voltage of th node as while treating the adjacent myelinated section as a classical Ohmic resistor with resistance (where is intracellular resistance per unit length and is the length of myelinated section) the current between nodes and can be written as
Considering the conservation of current at th node implies that the total transmembrane current into that node can be written as
where is the radius of the axon. Bressloff [26] extracts term from (14) and constructs equation for the potential at node as
where
Here , , and is the length of the node of Ranvier. The parameter (16) governs the saltatory conduction velocity between adjacent nodes of Ranvier under the simplifying assumptions done above and is a reasonable starting point. It is worth noting that the philosophy behind constructing eq. (15) is one of the typical approaches how the HH equations (which in its ODE form describes signal evolution in time at a fixed spatial point on axon) are used to describe travelling AP signal – meaning that one constructs a pseudo-numerical scheme where adjacent nodes interact and can propagate the signal that way along the axon.
We have to emphasize that what follows, is not the direct one-to-one adoption of the saltatory conduction mechanism as proposed in [26] as we are not only considering the electrical effects but also other influences from potentially relevant interactions (like cytoskeleton, for example) and as such, our approach is partly phenomenological. It should be stressed, that we opt to separate myelin geometry along the axon and perpendicular to the axon as theoretically easily observable and propose that the signal propagation velocity from node of Ranvier to node of Ranvier is governed by more than only membrane capacitor dynamics. As we know the actual propagation velocity from experimental observations (see, for example, [31] and references therein). Knowing AP velocity allows accounting of the influence of the membrane capacitor dynamics (which can be estimated from the myelination geometry and dielectric properties of the myelin). Then it is possible to estimate other potential influences, like fast enough chemical processes or interactions with the internal structures of the axon to quantify these and hopefully provide a more detailed insight towards functioning and signal propagation in an actual living nerve cell.
4 Model for AP propagation on myelinated axon based on Lieberstein model
Taking the elementary form of Maxwell equations combined with the ideas proposed by Lieberstein in eqs. (8) and (9) as a starting point, we proceed to modify these governing equations to include the effect of myelination on the AP signal propagation. The simplified geometry of a myelinated axon is shown in Fig. 5.
4.1 Phenomena and descriptions
When we modify the Lieberstein model [23] to account for the effect of myelination on a nerve fibre, we consider the following phenomena:
1.
The velocity of the AP depends on the ratio of lengths between the myelin sheath and the node of Ranvier (so-called ‘-ratio’ below);
2.
The thickness of the myelin sheath affects the velocity of the AP signal (the so-called -ratio) and could be taken into account indirectly through the capacitance variations (included in parameter below);
3.
The dominant mechanism through which the AP signal velocity in myelinated nerve fibre is increased is the so-called saltatory conduction hypothesis [26];
4.
The model equation should be reduced back to the basic model when the myelination approaches to zero (i.e., unmyelinated axon).
Let us take Lieberstein eqs. (8) and (9), introducing parameters and characterizing the AP propagation velocity increase from saltatory conduction [26] and other relevant mechanisms. Note that Bressloff [26] has used parameter that modulates the signal dynamics across the membrane. Here, inspired by Bressloff [26], we have introduced two parameters: and . The equations describing the propagation of the AP are written in the form:
| Table 3: AP velocity against -ratio. | |||||||
| : | 0 | 25 | 50 | 75 | 100 | 125 | 150 |
| µ | 1 | 5.3 | 7.5 | 9 | 10.5 | 11.7 | 12.8 |
| µ | 1.9 | 9.9 | 13.9 | 17 | 19.6 | 21.8 | 23.9 |
| µ | 2.5 | 13 | 18.2 | 22.3 | 25.7 | 28.6 | 31.4 |
| µ | 3.2 | 16.3 | 22.8 | 27.8 | 32 | 35.8 | 39.2 |
| µ | 3.8 | 19 | 26.7 | 32.7 | 37.6 | 42 | 46 |
| : | 175 | 200 | 225 | 250 | 275 | 300 | 325 |
| µ | 13.8 | 14.8 | 15.7 | 16.5 | 17.3 | 18 | 18.8 |
| µ | 25.9 | 27.6 | 29.3 | 30.9 | 32.4 | 33.8 | 35.2 |
| µ | 33.9 | 36.2 | 38.4 | 40.5 | 42.4 | 44.3 | 46.1 |
| µ | 42.3 | 45.2 | 47.9 | 50.5 | 53 | 55.3 | 57.7 |
| µ | 49.7 | 53 | 56.3 | 59.3 | 62.2 | 65 | 67.7 |
In eq. (17) parameter (describing -ratio) describes the average length of the myelinated section divided by the average length of the node of Ranvier (see Fig. 5). It affects the quantity which is the current along the axis of the axon. Parameter is a phenomenological coefficient which determines conduction velocity between adjacent nodes of Ranvier (generalised from eq. (16)). Here parameter includes myelin geometry perpendicular to the axon (related to -ratio). As noted earlier, parameter is not similar to parameter (see eq. (16)) and is a generalised quantity. We assume parameter to be between 0 and 1: if it is equal to 0, then the current can not propagate between the adjacent nodes (nodes of Ranvier are isolated from each other) and if it is equal to 1, then the myelinated sections are almost perfect conductors (i.e., the system obeys Ohm’s Law) and adjacent nodes react to the changes in a given node almost immediately. Parameter could be considered like a generalised dimensionless relative velocity potentially containing all the physical effects (not only resistance but also diffusive, capacitative, inductance effects, the influence of cytoskeleton or even damage or pathological effects, etc) as it is chosen here. One should also stress that in the following example, we take for the sake of simplicity, however, it would be logical that as the length of myelinated sections increases then at some point the parameter would start to decrease. Using the same parameters as in the classical paper by Hodgkin and Huxley [1], a numerical example using parameter value (assuming 1 µ nodes of Ranvier, 50 µ length for myelinated sections and taking for the numerical example) follows (see Fig. 6). Choosing (which is an upper limit case) means that we are making here an assumption that saltatory conduction velocity is much faster than signal propagation velocity in an unmyelinated axon. We have to emphasize here that the “unit cell" we are modelling in the computational node contains both the node of Ranvier and the myelinated section adjacent to it, i.e., the parameters are a mix of “pure” Lieberstein model for unmyelinated axon and myelination effects like in [26].
One can note that in time (at a fixed spatial node, Fig. 6) the solutions look practically identical. However, it must be noted that in space the faster (in myelinated axon) AP is shaped significantly wider, as the dynamics of the ion channels are still the same as before, but as the wavefront propagates faster, its dominant wavelength in space is longer.
The AP propagation velocity as a function of the -ratio at a fixed axon radius ( µ, µ, µ, µ, and µ) is depicted in Fig. 7. The case corresponds to an unmyelinated axon. It should be noted that here we are assuming that the thickness of the myelin (related to g-ratio, which is included through coefficient (16)) is the same even if -ratio is varied. The upper limit case () for the conduction velocity is used. It should be noted that the dependence of the AP velocity on the axon radius is still there even in the myelinated case. The larger the radius of the axon, the faster the velocity of the AP increases as the -ratio is increasing.
4.2 On AP energy in myelinated axon
When analysing the propagation of the AP, several forms of energy could be considered: capacitative, heat, inductive and even chemical. What follows is a short discussion on the AP energy based on our studies in addition to the previous analysis [21, 38, 37, 60, 61].
As a first rough estimate of just one part for the energy of AP, we can take it as the energy contained in an electric field on a capacitor. Taking the simplified axon geometry as a cylinder with a charge difference across the membrane one can make a naive simplification that the capacitance of such a cylinder segment is the same as a classical plate capacitor
where is the electric constant or vacuum permittivity, is the dielectric constant (roughly 7 for lipid bi-layer), is area (for cylinder wall where is the radius and is the length) and is the thickness of insulating layer between the plates. As an example, for a rough estimate let us take µ, and . Inserting the noted values into the (22), we get capacitance of roughly (per of axon). However, it must be noted, that the value of the dielectric constant can vary across protein internal structure and its changes [62].
The energy stored in the electric field in a charged parallel plate capacitor can be expressed as
where is capacitance and is potential difference. Assuming , we get a rough estimate of the energy as per cm.
This is a rough estimate for an unmyelinated axon (i.e., the node of Ranvier) because any geometry variations or possible changes of the dielectric constant during AP generation are neglected. Also, it must be noted that the electric field is quite significant in the unmyelinated sections (approximately potential difference across only roughly distance yields an electric field of ).
Now if we add into our consideration the myelination it seems reasonable to assume that energy must be conserved, i.e., when the AP signal is “jumping" between the nodes of Ranvier the energy of the pulse is roughly the same per cm as we estimated for the node of Ranvier where the main difference is that the thickness of the insulating layer is much larger. For example, let us take µ. In such a case the capacitance estimate drops to roughly per cm if all other parameters remain the same and the average electric field in the insulating layer drops to about for a myelination thickness.
However, one should note that in the myelinated case, the signal is substantially faster and as a result the AP is spread over a significantly longer section of the axon in the myelinated case. One can argue that if considering the “total" AP signal energy (over its whole duration, at a fixed spatial point or at a fixed time over its whole spatial distribution) it could be generalised (drawing inspiration from Poynting theorem in 1D case in classical electrodynamics [63]) to consider the integral instead of just in (23). Considering this idea we can combine eqs. (22) and (23) replacing the with its integral over half-space can write
as a rough estimate of the total energy of the electric field for AP in space or in time, respectively. Note that we have assumed here that only a single AP pulse exists in the half-space or within the given timeframe for eq. (24). It should be emphasized, that unlike classical capacitor energy expression (23) that yields estimate in joules the proposed generalised expressions (24) yield estimates on joule meters and seconds, respectively.
On the other hand, considering only the question of capacitor energy is only one aspect of the AP energy. One could ask how the inclusion of inductivity affects the calculation of AP energy? As a first approximation, one could be tempted to consider the energy in transmission line (1), (2) including inductivity (see Fig. 8 for a scheme of a unit cell) as opposed to just capacitor energy as it is usually done in studies based on the HH equation. However, authors of the present paper remain somewhat sceptical towards such an approach as it is not entirely clear if considering magnetic field energy on such a small scale in the electrically conductive environment inside an axon would yield a reasonable estimate, even if we insist on including the noted inductivity. After all, the model proposed here does not include the influence of the cytoskeleton explicitly (although its influence on the AP propagation is taken into account indirectly through parameter ).
Another possibility is to approach the question of the energy of AP from the perspective of power. It is known that power is , where is the current in a circuit and is the potential difference. Considering the power we could separate it into two generalised components, first, across the membrane at a fixed spatial point where is the current across the membrane (taken similar to the HH equations here) and is the AP amplitude at the given spatial point at given time (about 100 at the peak) and second, along the axon at the fixed time where is current along the axis of the axon and is the potential gradient along the axis of the axon. The total power of the electrical pulse would be then .
5 Discussion and summary
We started with the elementary form of Maxwell equations (1), (2) and drawing inspiration from the classical HH paper [1] used the total membrane current density (6) to derive the model equations for unmyelinated axon similar to the Lieberstein model [23]. While Lieberstein opted to go a step further by moving into a moving frame of reference we opted to stay with the form which is closer to the Maxwell equations for a transmission line eqs. (8), (9) and used the work of Lieberstein as a source of inspiration for handling the question of inductance in the context of signal propagation along the nerve axon. Using the parameters from the literature we investigated briefly the solutions of the noted model for the unmyelinated axon and demonstrated that the behaviour of the solutions is in the physiologically plausible range and the key characteristics of the nervous signalling are fulfilled. These are: (i) the annihilation of AP signals during a head-on collision, (ii) the existence of activation threshold, and (iii) the refraction period after signal passing. In the parameter range considered, we have calculated the AP signal propagation velocity from 0.5 up to about 4.4 for the unmyelinated axon. The key difference between the classical HH equations and the Lieberstein-inspired model used here is that the mechanism for signal propagation along the axon emerges like a wave as a consequence of opting to keep the inductivity . While, indeed, there exist variations of the classical HH equations which support AP signal propagation where the equations are written in the form of PDE instead of the usual ODE form (which describes signal evolution in time at a fixed spatial point). In these, normally, the potential-gradient-type member responsible for propagating the signal along the axon is not as clearly defined from the viewpoint of physics (usually some kind of abstracted diffusion-type process is used). This is essential later in the paper for the purpose of clearer physical interpretation as we make use of the model based on the elementary form of Maxwell equations which is further modified to include the influence of myelination on the signal propagating along the axon.
We briefly explained the saltatory conduction mechanism hypothesis as we prepared to modify the governing equations to take into account the effect of myelination on the propagation of the AP along the axon. The model (8), (9) based on Maxwell equations for a transmission line is then modified to include the influence of myelination. In addition to the g-ratio normally considered in the earlier studies (taken into account indirectly through coefficient ) which takes into account the myelination geometry perpendicular to the axon we introduce the so-called “myelination-ratio" or -ratio which describes the influence of myelin distribution on the signal propagation in the direction of the axis of the axon. The numerical example (see Fig. 7), using parameters from the earlier studies, demonstrates physiologically plausible behaviour for the model. The equations are reduced to the description of the unmyelinated axon if the length of the myelinated sections along the axon is taken as zero. Under the considered parameter combinations we can observe the AP propagation velocities up to 67.7 (the signal propagation velocity range for myelinated axons is given as roughly 10 to 100 in the earlier studies [29, 30]).
It is important to emphasize that the proposed continuum-based model is philosophically similar to how the transmission line equations are composed. The ‘unit-cell’ in the context of the myelinated axon in the model is composed of the node of Ranvier and the myelinated section next to it. This is opposed to the alternative approach where the processes in nodes of Ranvier are described by the classical HH equations while myelinated sections are handled separately either through some numerical scheme or by an alternative model coupled with the HH equations in the node of Ranvier through some mechanism. Having a relatively simple pair of PDEs which are connected to the Maxwell equations for anything involving the movement of charges in an environment could be considered superior to investigating causal connections and making predictions than something that is not as clearly connected to the basic physical considerations.
Finally, we briefly discuss the question of the energy of the AP pulse in both myelinated and unmyelinated axons, making an argument that while the normally used capacitor-energy-based energy estimate is a reasonable starting point, something more could be needed. Speculating that perhaps instead of considering capacitor energy as the energy of an AP signal it could be better to consider the power of the signal (either in space, at a fixed time or in time, at a fixed spatial point).
Acknowledgements
This research was supported by the Estonian Research Council (PRG 1227). Jüri Engelbrecht acknowledges the support from the Estonian Academy of Sciences. The authors thank Prof. Antoine Jerusalem from the University of Oxford for helpful comments during the preparation of the present paper. The authors thank the anonymous reviewers for numerous suggestions for improvements and helpful comments.
Author contributions Kert Tamm contributed by writing manuscript text and composing figures, performing numerical simulations and analysis of the results. Tanel Peets contributed analysis for comparing the modified Lieberstein equation against the classical Hodgkin-Huxley equations. All other work (like novel ideas presented, physical interpretation of the results, proofreading, etc) is a collaborative effort equally contributed by all authors.